About this document

The following pages are duplications of the section “New Publications”, found in each issue of TEX and TUG NEWS, the quarterly newsletter of the TEX Users Group.

The listings are augmented by margin notes indicating reviews which have subsequently appeared in TUGboat, TUG’s quarterly journal.

At the end are two pages listing books which had been reviewed in TUGboat before TTN began, as well as a listing of books on TEX in Russian (which originally appeared in TTN 3,4:5.

Updates of this file (ttnpubs) are posted to CTAN on a regular basis:

                \verb|tex-archive/digests/ttn}|

New Publications (TTN 1,1)

Christina Thiele Editor, TTN Claudio Beccari. LaTEX — Guida a un sistema di editoria elettronica. Milano: TUGboat 13, 2.

Editore Ulrico Hoepli Milano, 1991. Pp. ix + 399. ISBN 88-203-1931-4.

• The book is the first and only one in Italian at this moment. Although it contains an introductory chapter for the beginner, its main purpose is to give the reader advanced information on LaTEX (and partially on TEX) in order to prepare sophisticated macros, style files, and to make clever use of fonts.

Arvind Borde. TEX by Example. New York: Academic Press, 1991. Pp. 169 TUGboat 13, 4. (paper), US$19.95. ISBN 0-12-117650-9.

• This book teaches entirely by example. Input and output are shown sideby-side, allowing users to spot quickly how to get symbols, how to obtain italics, how to use bold type, how to start new paragraphs, and so on. The alphabetic appendix is both an index and a glossary, which systematically lists all commonly used features of TEX.

Victor Eijkhout. TEX by Topic: A TEXnician’s Reference. Wokingham, UK: TUGboat 13, 2. Addison-Wesley, 1992. Pp. viii + 307 (paper), US$30.00. ISBN 0-201-56882-9.

• TEX by Topic is a systematic reference manual designed as a companion to tutorial guides about TEX.

Not for the novice, it is meant for users with a basic understanding of TEX who have specific questions, or who are keen to explore the full potential of this unique computer typesetting system.

Jane Hahn. LaTEX for Everyone. Mill Valley: Personal TEX Inc., 1991. Pp. xi TUGboat 13, 2. + 346 (paper), US$19.95. ISBN 0-9631044-0-3.

• LaTEX for Everyone is a tutorial as well as a reference guide. Written with the beginning user in mind, it is designed to be easy to read and understand, and contains examples of all the basic commands. The experienced LaTEX user, however, will also find this book to be a valuable reference tool.

Raymond Seroul and Silvio Levy. A Beginner’s Book of TEX. New York: TUGboat 13, 1. Springer Verlag, 1991. Pp. 283 (paper), US$29.95, DM58,– . ISBN 0-387-97562-4.

• Seroul and Levy’s book is an excellent beginner’s text. It is meant for careful reading, in which case it gives a surprising amount of detail. The book also has a reference section. This is an expanded translation of Seroul’s book, Le petit Livre de TEX (Paris: InterEditions, 1989).

André Heck, ed. Desktop Publishing in Astronomy and Space Sciences. Sing- TUGboat 13, 4. apore: World Scientific Publishing Co. Pte. Ltd., 1992. Pp. xi + 240 ISBN 981-02-0915-0.

• The volume gathers together the communcations presented at the colloquium on Desktop Publishing in Astronomy and Space Sciences, held in Strasbourg, France, 1–3 October 1991, at the Louis Pasteur University.

Wynter Snow. TEX for the Beginner. Reading, Mass.: Addison-Wesley, 1992. TUGboat 13, 4. Pp. xii + 377 + 23 (index) (paper), US$29.25. ISBN 0-201-54799-6.

• This book is a carefully paced, tutorial introduction for people first learning the [TEX] system. Special emphasis is given to what can go wrong, and how to fix it when it does. LaTEX notes are provided for use with a set of macros.

Latest issues of other TEX newsletters

Les Cahiers GUTenberg: no. 12, décembre 1991, 85pp. ISSN 1140-9304. Official publication of GUTenberg, Groupe francophone des Utilisateurs de TEX.

MAPS: Issue 92.1, May 1992, 162pp. Official publication of NTG, Nederlandstalige TEX Gebruikersgroep (MAPS = “Minutes and APendiceS”). Two issues per year. Contact: ntg@hearn.bitnet.

TEXbulletin: no. 1, 1992, 48pp. Official newsletter (in Czech and Slovak) of CsTUG. Previous 4 issues in 1991 appeared without numbering. Four issues per year. Contacts: Karel Horak horakk@csearn.bitnet, Jiří Veselý

jvesely@cspguk11.bitnet, or Jiří Zlatuska zlatuska@cspuni12.bitnet.

TEXline: issue 14, February 1992, 28pp. Contact Malcolm Clark for details

malcolmc@sun.pcl.ac.uk.

Die TEXnische Komödie: vol. 4, no. 1, mai 1992, 68pp. Official newsletter of DANTE, Deutschsprachige Anwendervereinigung TEX e.V. Four issues per

year. Contact: dante@dhdurz1.bitnet or dante@vm.urz.uni-
heidelberg.de.

Note: For more information on the various user groups, consult pp. 119–125 in TUG’s 1991 Resource Directory.

Jiří Zlatuška, ed. EuroTEX 92, Proceedings of the 7th European TEX Conference. Brno: Masaryk University and Prague: Czechoslovak TEX Users Group (CsTUG), 1992. ISBN 80-210-0480-0.

• Copies can be ordered directly from CsTUG for the price of DM 30.– plus postage (or the equivalent in some other currency) at the following address: Československé sdružení uZivatelů TEXu (CsTUG), Matematický ústav UK, Sokolovská 83, CS-186 00 Praha 8, Czechoslovakia. They may also be available through your local TEX users group.^{1}

Stephan von Bechtolsheim. TEX in Practice. New York: Springer Verlag,

{TUGboat 15, 1; \atop 15, 2.}

Oct. 1992. 4-volume set (hardcover), US$169.00. ISBN 0-387-97296-X. Volumes also available individually, US$49.00 each:

Vol. 1: Basics. Pp. 359, 9 illus. ISBN 0-387-97595-0.

Vol. 2: Paragraphs, Math, and Fonts. Pp. 384. ISBN 0-387-97596-9. Vol. 3: Tokens, Macros. Pp. 544, 22 illus. ISBN 0-387-97597-7. Vol. 4: Output Routines, Tables. Pp. 300. ISBN 0-387-97598-5.

• “A recent surge of good TEX implementations for PCs has put TEX on the disks of many people including writers, designers, desktop publishers, and engineers. With such increased interest in TEX, there is a need for good TEX books. TEX in Practice is the ideal reference and guide for the TEX community. The four-volume set is written by an acknowledged expert in the field and addresses the needs of the TEX novice to the more experienced ‘TEXpert.’ The book provides step-by-step introduction to the various functions of TEX with many relevant examples.”

Stefan Schwartz and Rudolf Potuček. LaTEX — Satzkunst statt DTP. [“LaTEX — The Art of Typography rather than DTP”]. (Special issue in the series CHIP-Anwender-Praxis.) 1992. DM 34,– . ISBN 3-8023-1178-7.

• Can be obtained from: Vogel Verlag, Postfach 67 40, D-8700 Würzburg, Germany.

Latest issues of other TEX newsletters

Les Cahiers GUTenberg: no. 13 juin 1992, 90pp. ISSN 1140-9304.

1The Table of Contents of the Proceedings appears on p. 28 of this issue, and a short report on the conference by Alan Hoenig appears on pp. 22–23.

Rames Abdelhamid. Das Vieweg LaTEX-Buch: Eine praxisorientierte Einführung. Braunschweig: Vieweg-Verlag, 1992. About 200pp. ISBN 3-528-05145-0.

• This book provides a well-structured and readable introduction to LaTEX in combination with the german.sty option.

George Gratzer. Math into TEX. New York: Springer Verlag. Pp. 187 (paper). TUGboat 15, 1. Approx. $34.00 US. ISBN 0-8176-3637-4. [Avail. Dec. 1992]

• “This book is for the mathematician, engineer, or scientist who wants to write and typeset articles with mathematical formulas but who does not want to spend a great deal of time learning how to do it. It assumes little familiarity with TEX or LaTEX.” [From promotional material]

Piotr Wyrostek, trans. LaTEX: System przygotowywania dokumentöw. Przewodnik ürytkownika i prodrëcznik. Polish translation of Lamport’s LaTEX manual. Krakow: Ariel. Price: 91000 zl. ISBN 83-900460-1-6.

Michaela Lichá and Oldřich Ulrych. AMS-TEX verze 2.1. Prague: Czechoslovak TEX Users Group, 1992. Pp. 80 [In Czech]

• “This publication provides a short (but complete) tour through AMS-TEX 2.1 together with amsppt.sty 2.1. Almost all features are demonstrated with a lot of examples. The book is organized in such a way that it is useful for novices (basic textbook), and for advanced users (quick reference quide) as well. Contains a table of all symbols.” [Info. from authors]

Latest issues of other TEX newsletters

Baskerville: vol. 2, no. 1, March 1992, 12pp. Official newsletter of the UK TEX

Users’ Group. Die TEXnische Komödie: vol. 4, no. 2, August 1992, 64pp. Official newsletter of

DANTE, Deutschsprachige Anwendervereinigung TEX e.V. MAPS: Issue 92.2, 1992, 176pp. Official publication of NTG, Nederlandstalige

TEX Gebruikersgroep (MAPS = “Minutes and APendiceS”). TEXbulletin: no. 2, 1992, 48pp; no. 3, 1992, 64pp. Official newsletter of CsTUG.^{2} TUGboat: 13, no. 3, 1992, 162pp. Proceedings of TUG’92 meeting in Portland,

Oregon.

2Re: TTN 1,3 notice on the Prague Proceedings: Jiří Zlatuska of CsTUG asks that orders for the publication include a shipping and handling charge: $5 for surface mail, $10 for airmail in addition to the 30,– DM cost. –Ed.

Donald E. Knuth. Literate Programming. (CSLI Lecture Notes 27.) Stanford: Center for the Study of Language and Information, 1992. Pp. xiii + 368 £19.95 (UK). ISBN 0-937073-80-6 (paper), ISBN 0-937073-81-4 (cloth).

• ‘Literate Programming’ (using the WEB system) was developed by Knuth and used to write TEX and its documentation. This book is a collection of previously published articles and extracts from books, together with additional references and some other new material. In particular, it contains the papers “The Errors of TEX” and the “Error Log for TEX, 1978–91”.

Michael Vulis. Modern TEX and Its Applications. Boca Raton: CRC Publish- TUGboat 14, 1. ers, 1992. Pp. 275, includes diskette. US$32.95 ($39.95 outside the US). ISBN 0-8493-4431-X (paper).

• “A guide to plain TEX (and VTEX), intended for scientists, etc., and support staff who prepare technical texts and documents. It covers all basic topics for understanding the TEX user environment, including fonts and characters, formatting, math mode, macros, terminal and file operations, tables, foreign language capabilities (including Cyrillic), indices, references, tables of contents, scalable typefaces, and graphics. A PC-compatible disk containing examples, extra typefaces, and a ready-to-run restricted [!] version of TEX is included.” [adapted from promotional material]

Antoni Diller. LaTEX Line by Line: Tips and Techniques for Document Pro- TUGboat 14, 2. cessing. (Software Engineering.) Chichester: John Wiley and Sons, 1992. Pp. 200 (paper). US$29.95. ISBN 0-471-93471-2.

• “A handbook to the LaTEX text processing system, this easy-to-follow guide teaches all users how to produce a wide variety of documents — from business letters to technical reports. The book offers ready-to-use templates for producing articles, reports, letters and more; explains how to quickly generate tables of contents, indices and bibliographies; and includes many practical examples, an appendix of mathematical symbols and a complete glossary of typesetting terms.” [adapted from promotional material]

G. Grätzer. Math into TEX: A Simplified Introduction Using AMS-LaTEX.

Basel: Birkhäuser, 1992. SFr 52.–, DM 57.–. ISBN 3-7643-3637-4. Note: This item was already listed in the last issue. This entry adds the bibliographic data of the co-publisher.

Latest issues of other TEX newsletters

Die TEXnische Komödie: vol. 4, no. 3, November 1992, 56pp (DANTE).

Malcolm Clark. A plain TEX Primer. Oxford: Oxford University Press, 1992. Pp. 481. ISBN 0-19-853784-0 (cloth), ISBN 0-19-853724-7 (paper).

• “This primer introduces plain TEX, and aims to help the reader get started. It explains why TEX approaches its subject in the way it does, and provides the context into which it fits. Another thread is the emphasis placed on document structure. Stress is also laid on practical work. The book is aimed at scientists and researchers preparing their own books and papers, or technical typists used to the conventions and jargon of their field.” [adapted from the publisher’s description]

Stanley A. Sawyer & Steven Krantz. A TEX Primer for Scientists. Boca Raton: CRC Publishers, 1993. Pp. 175. US$29.95 ($36.00 outside USA). ISBN 0-8493-7159-7.

• “This concise, straightforward text covers all basic topics of equation typesetting, and some advanced topics: how to automatically generate a table of contents, and indices; how to make text flow around a graphic; how to add more fonts; how to change fonts in math mode (for titles, etc.). In an appendix differences between TEX, LaTEX and AMS-TEX are discussed.” [adapted from promotional material]

Hans Ibrus. Väike Eesti TEXiraamat. [ The Small Estonia TEXbook ]. Tartumaa, Estonia: a/S “KompuT” kirjastus, 1992. Pp. 144.

• Preface by Enn Saar. [Further description on p. 18.] Appendix A is a translation into Estonian of First Grade TEX, by Arthur Samuel.

Gordon Müllejans. METAFONT. Eine Referenz. Bonn: Addison-Wesley, 1992. Pp. 176. DM 39,90. ISBN 3-89319-363-4.

• This is neither a tutorial nor a textbook on . It is a reference tool, and consists mainly of listings with short descriptions, in particular of the basic commands, the macros of cmbase, and the parameters used to define the CM fonts.

Latest issues of other TEX newsletters

Die TEXnische Komödie: vol. 4, no. 4, Dezember 1992, 56pp.

La Lettre GUTenberg: numéro 0, février 1993, 16pp.

Helmut Kopka and Patrick W. Daly. A Guide to LaTEX: Document Prepa- TUGboat 14, 2. ration for Beginners and Advanced Users. Addison-Wesley (UK) Ltd., 1993. Pp. 448. US$34.50. ISBN 0-201-56889-6.

• This book is based on two well-known German books by Helmut Kopka: LaTEX — Eine Einführung, and LaTEX — Erweiterungsmöglichkeiten.

Francesc Rosselló Llompart. L’ABC del TEX. Palma de Mallorca: UIB-DMI.

Departament de Ciències Matemàtiques i Informàtica, Universitat de les Illes Balears, 1991.

Reinhard Wonneberger. LaTEX. 3., durchgesehene und erweiterte Auflage.

(Addison-Wesley Kompaktführer). Bonn, etc.: Addison-Wesley, 1993. Pp. xvi + 166. DM 29,80. ISBN 3-89319-589-0.

• This is the third (expanded) edition of a concise German LaTEX reference (previous editions 1987 and 1988). The main addition is a section on german.sty, which is widely used in German-speaking countries.

Electronic Publications

Hubert Partl, Elis. Schlegl and Irene Hyna. Una Descripción de LaTEX.

• Spanish translation by Tomas Bautista of well-known German introduction to LaTEX, LaTEX-Kurzbeschreibung. Includes material on the spanish style option. Available in Spain from ftp.eunet.es (also known as goya.

uu.es) in the directory info/unix/text/TeX\spanish/tutorial/, as well

as from the main TEX archives.

Latest issues of other TEX newsletters

Les Cahiers GUTenberg: No. 14: EuroTEX’92: Proceedings of the 7th European TEX Conference. Prague, Sept. 14–18, 1992. Pp. vi + 330. No. 15, avril 1993, 61pp. ISSN 1140-9304.

GUST: Zyszyt 1 (1993). [March issue 35 + 7; A4 format.]

• First issue of the newsletter GUST from the Polish user group Grupa Użytkowników Systemu TEX. Editors are: Włodek Bzyl and Tomasz Przechlewski. Address is: Instytut Matematyki UG, Wita Stwosza 56, 80-952

Gdańsk. E-mail: matwb@halina.univ.gda.pl.

MAPs: Issue 93.1, 1993, 233pp. [Dutch-speaking user group, NTG publication].

Arvind Borde and Tomas Rokicki: A User’s Guide for TEXHelp; The On-Line TEX Handbook. Academic Press, 1993. Pp. v + 28. ISBN 0-12-1176040-1.

• This small booklet documents the TEXHelp program for IBM PC and compatible microcomputers.

Arvind Borde: Mathematical TEX by Example. Academic Press, 1993. Pp.

356. US$19.95. ISBN 0-12-117645-2.

• As the title suggests, this book is an introduction to TEX by means of examples. They almost entirely consist of pairs of pages: one contains some piece of TEX code (with explanations and comments added via footnotes), while the other shows the output produced by this code.

Nicholas J. Higham: Handbook of Writing for the Mathematical Sciences.

SIAM — Society for Industrial and Applied Mathematics, Philadelphia, 1992. Pp. xii + 241. US$21.50. ISBN 0-89871-314-5.

• This book is intended for all authors of technical papers (and books) or talks, and covers virtually all relevant issues. The author discusses questions of style as well as practical matters. The use of computers in the publishing process — including TEX, of course — is discussed in detail. An appendix contains a list of TEX and LaTEX symbols. Among the various useful materials are hints for Internet use and an extensive bibliography.

Christian Rolland: LaTEX guide pratique. Addison-Wesley France, July 1993. TUGboat 15,2 Pp. 280. approx. 190 FF. ISBN 2-87908-025-0. (Distributed by Bordas, Bordas code 009808.)

Newsletters

GUST: Zeszyt 2 (1993), 42pp. Official newsletter of the Polish User Group, Grupa Użytkowników Systemu TEX.

La Lettre GUTenberg: no. 1, juillet 1993, 16pp.

Die TEXnische Komödie: 5. Jahrgang, Heft 1, Mai 1993, 72pp; Heft 2, Sep. 1993.

TUGboat: vol. 14, no. 1, April 1993, 92pp; vol. 14, no. 2, July 1993, 64pp.

Michael Doob. TEX: Starting from 1 [“square one”]. Springer Verlag, 1993. Pp. viii + 122 (paper). US$25.00 ISBN 0-387-56441-1, ISBN 3-540-56441-1.

• “A book designed for the complete newcomer to TEX. [It] begins with simple exercises on typesetting text and slowly advances into more complex problems such as different types of mathematical constructions and tables. The book presents a comprehensive overview of TEX with various tables included for quick reference. Readers will learn how to write standard papers of modest complexity with TEX and obtain the fundamental basis for learning more complex typesetting functions.” [from promotion material]

Francisco Ortegón Gallego. LaTEX, primeros pasos [LaTEX, First Steps]. Barcelona: Manuales de Informática Masson, Masson, S.A., 1992. Pp. xix + 168. 2,450 pesetas (approx. US$19). ISBN 84-311-0649-2.

• This book describes the main features of the LaTEX format. It is intended for beginners, but it can be interesting for initiated users, too. The book is divided into six chapters (here called lessons) and three appendices. Lesson one introduces many topics by discussing a complete example of a source file. [There exists an erratum file at —ftp.math.utah.edu— in —pub/tex/errata— named —p-ortegon-gallego.ltx—.] [from a description by the author]

Eitan M. Gurari. Writing with TEX. New York: McGraw-Hill, 1994. Pp. xiv + 249. US$29.95. ISBN 0-07-025207-6.

Eitan M. Gurari. TEX and LaTEX: Drawing and Literate Programming. New York: McGraw-Hill, 1994. Pp. xiv + 310. US$34.95. ISBN 0-07-025208-4 (book), ISBN 0-07-911616-7 (book + diskette).

Peter Scherber, ed. Offizin. Schriftenreihe zu TEX, LaTEX, und , Band 1. Published by DANTE e.V. Addison-Wesley Deutschland, 1994. ca. 190pp (hardcover). DM 49,90. ISBN 3-89319-396-0.

• This is intended as the first volume in a series of yearbooks published by DANTE e.V, the German-speaking TEX users group. It has been announced for a long time and is available now.

Lothar Schumann. Professioneller Buchsatz mit TEX: Lehrbuch für Anwender [Professional Book Typesetting with TEX: A Textbook for Users]. (Weiterbildung Informatik). München-Wien: R. Oldenbourg Verlag, 1989 and 1991. ca. DM 40. ISBN 3-486-21736-4 (1989: ISBN 3-486-21173-0).

• As a special feature this book contains its own (verbatim) source code as an appendix.

C. Vanoirbeek and G. Coray, eds. EP92. Proceedings of Electronic Publishing, 1992. (The Cambridge Series on Electronic Publishing). Cambridge University Press, 1992. Pp. x + 316 (hardcover). £35. ISBN 0-521-43277-4.

• This is the proceedings volume on EP92. Electronic Publishing ’92. International Conference on Electronic Publishing, Document Manipulation, and Typography (Swiss Federal Institute of Technology, Lausanne, Switzerland, April 7–10, 1992). Five of the twenty-one papers deal with topics closely connected with TEX (see Articles section, below).

Articles

Angelika Binding. “Einsatz von TEX beim Springer-Verlag.” Mitteilungen der Deutschen Mathematiker-Vereinigung 4 (1993), 19–22.

• This article (by a member of the German Springer branch) describes the use of TEX in the publishing process at Springer.

George Grätzer. “AMS-LaTEX.” Notices of the American Mathematical Society vol. 40, no. 2 (February 1993), 148–150.

George Grätzer. Advances in TEX Implementations. I. PostScript Fonts.” Notices of the American Mathematical Society: vol. 40, no. 7 (Sepember 1993), 834–838.

Suzanne M. Molnar. “A Tour of Graphical Typesetting on the Macintosh.” The UMAP Journal vol. 14, no. 2 (Summer 1993), 97–112.

• The topics discussed in this article include PICTEX and OzTEX.

Richard Morin. “Tangled up in TEX.” SunExpert Magazine vol. 4, no. 10 (October 1993), 42–46.

• A brief review of TEX support, including the Comprehensive TEX Archive Network (CTAN), the TEX Users Group, and several books on TEX.

Frank Mittelbach and Chris Rowley. “The Pursuit of Quality: How Can Automated Typesetting Achieve the Highest Standards of Craft Typography?” EP92. Proceedings of Electronic Publishing 1992 , 261–273.

Makoto Murata and Koichi Hayashi. “Formatter Hierarchy for Structured Documents.” EP92. Proceedings of Electronic Publishing 1992 , 77–94.

Gilbert B. Porter III and Emil V. Rainero. “Document Reconstruction: A System for Recovering Document Structure from Layout.” EP92. Proceedings of Electronic Publishing 1992 , 127–141.

Luigi Semenzato and Edward Wang. “Aleph: A Language for Typesetting.” EP92. Proceedings of Electronic Publishing 1992 : 65–75.

Richard Southall. “Presentation Rules and Rules of Composition in the Formatting of Complex Text.” EP92. Proceedings of Electronic Publishing 1992 , 275–290.

Newsletters

Baskerville: vol. 3, no. 2, December 1993, 24pp.

• The editor, Sebastian Rahtz, announces that Baskerville will appear six times a year.

MAPs: Issue 93.2, 1993, 242pp.

TUGboat: vol. 14, no. 3, October 1993, 212pp.

• This issue contains the proceedings of the Fourteenth Annual Meeting, Aston University, Birmingham, England, July 26–30, 1993.

TUGboat: vol. 14, no. 4, December 1993, 84pp.

[Note: There was a misprint in the last digit of the ISBN for Christian Rolland’s book, listed in TTN 2,4:3. The correct number is ISBN 2-87908-025-8. –PS] P.W. Abrahams, K. Berry and K. Hargreaves. TEX per l’impaziente

[TEX for the Impatient]. Milano: Addison-Wesley, 1991. Pp. 396. ISBN TUGboat 14,4 88-7192-022-8.

• This is a translation (by Gaia and Guido Franchi) of the well-known introduction, amplified by material relevant to the Italian language.

Michel Goossens, Frank Mittelbach and Alexander Samarin.

The LaTEX Companion. Addison-Wesley, 1993. Pp. 480 (paper). US$41.25. TUGboat 15,2 ISBN 0-201-54199-8.

• This book is based on LaTEX2e, the new version of LaTEX which is currently available in its betatest release. It explains tools and techniques that enhance the use of LaTEX and help format documents more quickly and more efficiently. Topics treated include customizing commands and enviroments, changing page layout, preparing indices and bibliographies, and the New Font Selection Scheme (NFSS), as well as using PostScript fonts and PostScript images. [adapted from promotional material]

Helmut Kopka. LaTEX — Einführung [LaTEX — Introduction]. Addison-Wes-ley, 1994.

• The new edition of Kopka’s well-know (German) books on LaTEX (also available in a partial English translation) includes LaTEX2e and will comprise three (instead of two) volumes. The first volume is already available.

Norman Walsh. Making TEX Work. O’Reilly and Associates, April 1994. (ap- TUGboat 15,2 prox.) 500pp. US$29.95. ISBN 1-56592-051-1.

• There are many books on TEX and LaTEX, but so far (it seems) no book has been devoted to the TEX system as a whole, i.e., all the software, tools, and macros that are available and can be used when working with TEX. This book tries to fill the gap and provides a comprehensive guide to the world of TEX. It discusses the various implementations of TEX on all common platforms and how to obtain them, presents the major macro packages, discusses the installation and the use of fonts (including PostScript and TrueType fonts), and the inclusion of graphics as well as the preparation of indices and bibliographies. It will certainly be a valuable source of information, in particular, for all those TEX users who lack net access.

Articles

George Grätzer. “Advances in TEX Implementations. II. Integrated Environments.” Notices of the American Mathematical Society, vol. 41, no. 2 (February 1994), 106–111 • “In this article I will review two integrated TEX implementations. Textures, the classical implementation on the Macintosh (in its modern variant, Lightning Textures), and the brand new PCTEX for Windows for IBM compatible computers. I shall also review Scientific Word, a TEX preprocessor.” [from the introduction]

Allyn Jackson. “What’s Happening: New Publication Takes the AMS in New Directions.” Notices of the AMS, vol. 41, no. 6 (July/August 1994), 626– 627

• This article presents the concepts of a new publication. The following is an excerpt from the production remarks: “Though the design is certainly more complex than the Society’s other TEX documents, TEX was chosen over a what-you-see-is-what-you-get desktop publishing program for its superior typographic quality. (A few small sections of the book were produced using a desktop publishing program but the output was manipulated and massaged by hand to make it look as good as the TEX output.)”

Tom Scavo, Yannis Haralambous, and Werenfried Spit. “TEX on the Macintosh.” Notices of the AMS, vol. 40, no. 10 (December 1993), 1353– 1360

• “There are currently four implementations of TEX available for the Macintosh family of computers: CMacTEX, DirectTEX, OzTEX, and Textures. The latter is a well-known commercial product which won’t be reviewed here; the others are either freeware or shareware and are the focus of this report.” [from the introduction]

Newsletters

Baskerville: vol. 4, no. 1, Feb. 1994, 24pp; vol. 4, no. 2, Apr. 1994, 24pp.

GUST: Zeszyt 3 (1994), 50pp. Official newsletter of the Polish TEX User Group.

La Lettre GUTenberg: no. 2, février 1994, 24pp [using Lucida-Bright]

Die TEXnische Komödie: 5. Jahrgang, Heft 3, Dezember 1993, 68pp.

Abass Alamnehe. eLaTEX: Document Preparation Guide [in Amharic]. EthiO Systems, 1993. Pp. xvi + 230.

• This is a manual to eLaTEX, an Ethiopian-based version of LaTEX.

Jonathan Grosvenor, Kay Morrison & Alexander Pim. The PostScript Font Handbook: A Directory of Type 1 Fonts. Revised edition. Wokingham: Addison-Wesley, 1992. Pp. ix + 425. US$41.50. ISBN 0-201-56893-4 (paper).

• For all those interested (or forced) to use PostScript fonts this directory provides a comprehensive collection of font samples from the font libraries of Adobe, Monotype, Linotype, and Agfa. Each font is represented by a full-page profile including the character set, a brief history, and hints for recommended usage. It includes an introduction on PostScript fonts.

Leslie Lamport. LaTEX: A Document Preparation System. 2nd edition. Addison-Wesley, 1994. Pp. 180. US$41.50. ISBN 0-201-52983-1 (paper).

• This new edition of the official LaTEX manual describes LaTEX2e. It is announced to be available soon.

Articles

Marvin S. Margolis. “MG Mathematical Graphics System.” Notices of the American Mathematical Society, vol. 41, no. 3 (March 1994), 200–201.

• “The MG Mathematical Graphics System creates and displays two- and three-dimensional mathematical graphics on an MS-DOS-based personal computer and produces high-quality PostScript output. The developers, R.B. Israel and R.A. Adams, […] primarily designed the relatively small program to assist authors to create graphs to include in TEX-typeset documents. It is available from MG Software, 4223 West Ninth Avenue, Vancouver, B.C., Canada V6R 2C6.” [from the introduction of this review]

Heinz Werntges. “Grafik-Import in LaTEX” [Importing graphics under LaTEX]. c’t 12/92, 252–258.

• This survey article (in the German computer magazine, c’t) discusses the various possibilities for including graphics into LaTEX documents. It has since been reprinted (slightly updated) in Die TEXnische Komödie, vol. 5, no. 2 (1993), 38–53.

Newsletters

Baskerville: Vol. 4, no. 3, June 1994, 24pp. Les Cahiers GUTenberg: No. 16, février 1994, 82pp. ISSN 1140-9304. [Thematic

issue: PSTricks et Seminar.] No. 17, mai 1994, 66pp.

La Lettre GUTenberg: No. 3, juin 1994, 24pp. [using Palatino]. MAPs: Issue 94.1, 1994, 178pp. [Dutch-speaking user group, NTG publication.] Die TEXnische Komödie: 5. Jahrgang, Heft 4/1993, April 1994, 60pp. Official

newsletter of DANTE e.V.

Luzia Dietsche and Joachim Lammarsch. LaTEX zum Loslegen — Ein Soforthelfer für den Alltag [LaTEX, ready to start — Instant help for every day].

Springer Verlag, Berlin-Heidelberg 1994. DM 39.00. ISBN3-540-56545-0.

• This is an introduction to LaTEX (2.09) based on courses which are held regularly by the authors.

Friedhelm Sowa. TEX, LaTEX und Graphik — ein Überblick über die Verfahren [TEX, LaTEX and graphics — a survey of the methods].

Springer Verlag, Berlin-Heidelberg 1994. DM 39.00. ISBN3-540-56468-3.

• This book (by the author of BM2FONT) is probably the first book (in any language) devoted to the inclusion of graphics into TEX documents. It treats macro packages producing pictures, the inclusion of bitmap graphics (BM2FONT) and vector graphics (hp2xx), and the inclusion of graphics by dvi drivers (in particular, PostScript by dvips).

Articles

George Grätzer. “Advances in TEX Implementations. III. A New Version of LaTEX, Finally.” Notices of the American Mathematical Society, vol. 41, no. 6 (July/August 1994), 611–615.

• This is a short introduction to, and a first review of, LaTEX2e.

Newsletters

Baskerville: Vol. 4, no. 4, Sept. 1994, 24pp.

MAPs: Issue 94.2, 1994, 232pp. [Dutch-speaking user group, NTG publication.]

Die TEXnische Komödie: 6. Jahrgang: Heft 1/1994 (Juli 1994), 68pp; and Heft 2/ 1994 (Sept. 1994), 60pp.

Zpravodaj [“Bulletin”]: vol. 1, 1994. [Official publication of the Czech/Slovak

CSTUG.]

Other books

Paul Abrahams. TEX for the Impatient. Reading, Mass.: Addison-Wesley, TUGboat 11, 4. 1990. 357 pp (paper).

• “TEX for the Impatient addresses the practical demands of busy people who have discovered the unmatched benefits of TEX, but whose time and interest in mastering all its intricacies may be limited … anyone … who needs a quick reference guide to TEX.”

Update: An Italian edition, TEX per l’impaziente, was published in 1991. TUGboat 14, 4.

Wolfgang Appelt. TEX für Fortgeschrittene. [TEX for the Advanced]. Bonn: TUGboat 11, 2. Addison-Wesley, 1988. ISBN 3-89319-115-1.

Arvind Borde. Mathematical TEX by Example. Academic Press, 1993. ISBN TUGboat 14, 1. 0-12-117645-2. Pp. 356.

David J. Buerger. LaTEX for Engineers and Scientists. McGraw-Hill, 1990. TUGboat 12, 2. ISBN 0-07-008845-4. Pp. 198 (paper).

Gianni Gilardi. Il TeX — Introduzione al linguaggio e complementi avanzati. TUGboat 15, 1. Zanichelli/Decibel, Bologna, 1993. ISBN 88-08-10860-0. Pp. 226 (cloth).

Helmut Kopka. LaTEX — Eine Einführung. 3rd rev. ed. Bonn: Addison-Wes- TUGboat 13, 1. ley, 1991. 375pp (hardcover).

Helmut Kopka. LaTEX — Erweiterungsmöglichkeiten mit einter Einführung in . 2nd rev. ed. Bonn: Addison-Wesley, 1991. 552pp. (hard- TUGboat 13, 1. cover).

Norbert Schwartz. Einführung in TEX. Bonn, Addison-Wesley, 1988. ISBN 3-925118-97-7.

Norbert Schwartz. Inleiding TEX. Addison-Wesley Europe, Amsterdam, 1990.

ISBN 90-6789-151-7.

Norbert Schwartz. Introduction to TEX. Reading, Mass.: Addison-Wesley, TUGboat 11, 2. 1990. 278 pp (paper). [Based on 2nd ed. of German original Einführung in TEX.]

Michael Urban. Premiers pas en LaTEX. Adaptation française de Éric Cornelis.

GUTenberg, 1991. 87pp (paper).

Russian books about TEX^{1}

1. Michael Spivak, Voskhititel’nii TEX. Moscow: Mir Publishers, 1993, 285pp. [Russian translation of The Joy of TEX by M. Spivak.]

The translator is Irina Makhovaya, the editor of the Russian version is Andrei Khodulev. This is the first book about TEX in Russian (7,000 copies).

2. Donald Knuth, Vsio pro TEX. Protvino: AO RdTEX, 1993, 575pp. [Russian translation of The TEXbook by Donald E. Knuth.]

The translator is Marina Lisina; the editors of the Russian version are Stanislav Klimenko and Sergei Sokolov. This book was published almost simultaneously with the Spivak book. The translation was very hard work and successful (10,000 copies).

3. Marat Evgrafov, Leonid Evgrafov, TEX: Rukovodstvo po naboru i redaktirovaniiu mathematicheskikh tekstov. Moscow: Phismathlit, 1993, 80pp.

The first Russian book about TEX that was written by Russians (7,000 copies).

4. Boris Shirokov, PCTEX. Prostoi sposob iziaschnogo oformleniia mathematicheskikh tekstov: uchebnoe posobie. Petrozavodsk: Petrozavodsk State University, 1994, 80pp.

This small manual talks about working with TEX on IBM PC compatible computers (not about PCTEX!) (300 copies).

5. Renat Zagretdinov et al., Izdatel’skaiia systema LaTEX. Kratkoe rukovodstvo. Kazan: Kazan State University, 1994, 94pp.

This manual is a short introduction to LaTEX 2.09, with some material added about the Russian version CyrTUG-emTEX; it also contains a glossary of LaTEX control sequences, etc. (1,000 copies).

6. Sergei L’vovskii, Nabor i verstka v pakete LaTEX. Moscow: Kosmosinform, 1994, 328pp.

The first complete description of LaTEX 2.09 in Russian (5,000 copies).

1NOTE: These books were listed in TTN 3,3:5.